# Learning Visual Proxy for Compositional Zero-Shot Learning

### v2

@@ Abstract
Compositional Zero-Shot Learning (CZSL) aims to recognize novel attribute-object compositions by leveraging knowledge from seen compositions. Existing methods align textual prototypes with visual features through Vision-Language Models (VLMs), but they face two key limitations: (1) modality gaps hinder the discrimination of semantically similar composition pairs, and (2) single-modal textual prototypes lack fine-grained visual cues, creating bottlenecks in VLM-based CZSL. In this paper, we introduce Visual Proxy Learning, a novel approach that facilitates the learning of distinct visual distributions, effectively reducing the modality gap and improving compositional generalization performance. Specifically, we initialize visual proxies for various attributes, objects, and their compositions using text representations. By optimizing the visual space, we capture fine-grained visual cues and guide the learning of more discriminative visual representations for attributes, objects and compositions. Furthermore, we propose an effective Cross-Modal Joint Learning (CMJL) strategy that imposes cross-modal constraints between the original text-image space and the fine-grained visual space. This approach not only boosts generalization for previously unseen composition pairs but also sharpens the discrimination of similar pairs, fostering more robust and precise learning. Extensive experiments demonstrate state-of-the-art performance in closed-world scenarios and competitive open-world results across four established CZSL benchmarks, validating the effectiveness of our approach in advancing compositional generalization.
@@ 1 Introduction
Figure 1: (a) The modality gap and coarse granularity of textual prototypes blur the inter-class boundaries in the visual space, causing classifiers to misclassify similar samples (e.g., slightly red unripe apples). (b) The visual proxy improves modality alignment, incorporates fine-grained image features, and makes the visual space more compact and separable, ensuring correct classification of similar samples.
@@
Thanks to the advancements in large pre-trained vision-language models like CLIP [31], recent research has made significant progress by leveraging CLIP and exploring various approaches to enhance its capabilities. These approaches include designing task-specific prompts [17, 26, 36, 37, 19, 3], decomposing compositional text features [18], and developing modal fusion methods [7] to narrow the modality gap. Such efforts have improved the alignment between compositional labels and image features, enabling CLIP to perform effectively in CZSL tasks. However, several challenges and inherent problems in the CZSL task still deserve attention.
Persistent Modality Gap. Despite efforts to narrow the modality gap between vision and language in existing methods, it persists as a fundamental challenge in CZSL. While prior work [16] demonstrates that the gap can be partially reduced, complete elimination remains unattainable, particularly under the current task setup where classification depends critically on top-1 cross-modal retrieval. In such scenarios, the distance between true cross-modal pairs may exceed that of false pairs, leading to misclassifications directly attributable to the modality gap (see Fig. 1(a)). This problem is further amplified by the fine-grained nature of compositional semantics, where subtle visual distinctions—such as texture variations between ”ripe” and ”unripe” apples—are semantically critical yet notoriously difficult to align with their textual counterparts.
Limited Discriminability of Textual Prototypes. Existing VLM-based CZSL methods learn textual prototypes to address label scarcity, yet these prototypes inherently lack the fine-grained visual details crucial for distinguishing similar compositions. While each visual category contains diverse image instances (e.g., apples in varying shapes or lighting conditions), textual prototypes are derived from a single compositional label, resulting in a semantic-visual asymmetry. This mismatch prevents prototypes from capturing subtle distinctions (e.g., ”ripe” vs. ”unripe” apples), leading to ambiguous decision boundaries in the visual space (Fig. 1(a)). As shown in prior studies [29], relying solely on textual prototypes fails to approach the optimal performance for visual tasks.
@@ 2 Related Work
Compositional Zero-Shot Learning. Previous methods for CZSL can be categorized into two strategies: (1) Uniform feature representation learning, which establishes relationships between compositions and primitives (attributes and objects), embedding both images and compositions into a shared space for unseen composition prediction using a single classifier [28, 24, 21, 22, 1, 13]. (2) Multi-branch feature learning, which employs parallel discriminative modules for attributes, objects, and compositions, emphasizing visual feature discriminability [25, 10, 12, 34, 33, 15]. Recently, with the advancement of pre-trained Vision-Language Models (VLMs) and their strong generalization abilities, VLMs have been increasingly applied to CZSL [26, 36, 19, 18]. For instance, Nayak et al. [26] used soft prompts to treat attributes and objects as learnable tokens, while Xu et al. [37] constructed a graph[35] of object-attribute compositions. Lu et al. [18] decomposed states and objects in language features and integrated them with image features. Huang et al. [7] created three recognition branches to model attributes, objects, and compositions, aligning branch-specific prompt representations with visual features. Li et al. [14] proposed a context-based, diversity-driven specificity learning framework, considering attribute specificity levels.
@@ 3.1 Problem Formulation
In CZSL, compositional labels consist of attributes and objects. Attributes are denoted as $a\in\mathcal{A}$, objects as $o\in\mathcal{O}$, and compositions as $c\in\mathcal{C}$, where $\mathcal{C}=\mathcal{A}\times\mathcal{O}$. The compositional label space is divided into seen compositions $\mathcal{C}_{s}$ and unseen compositions $\mathcal{C}_{u}$, with $\mathcal{C}_{s}\cap\mathcal{C}_{u}=\varnothing$. The training set is defined as $\mathcal{T}={(x,c),|,x\in\mathcal{X},c\in\mathcal{C}_{s}}$, where $\mathcal{X}$ represents the image space. During testing, the model must classify images from both seen and unseen compositions. CZSL can be categorized into closed-world and open-world settings based on the treatment of unseen compositions [20]. In the closed-world setting, unseen compositions $\mathcal{C}_{u}$ are provided as prior knowledge, and the prediction space is $\mathcal{C}_{pred}=\mathcal{C}_{s}\cup\mathcal{C}_{u}$. In contrast, in the open-world setting, unseen compositions include all possible attribute-object compositions, and the prediction space is $\mathcal{C}_{pred}=\mathcal{A}\times\mathcal{O}$.
@@ 4.1 Experiments Setting
Table 2: The experimental results for the closed-world setting on the VAW-CZSL dataset. The best performance is highlighted in bold.
@@
Table 3: Results of ablation experiments on dual-modal center learning. ’TP’/’VP’ represents the text prototypes and visual proxies.
@@
4.2 Main Results
We comprehensively evaluate our model under both closed-world and open-world settings, comparing it with recent CLIP-based CZSL methods from the past two years, including: CLIP[31], CoOp[39], CSP[26], DFSP[18], DLM[6], ProLT[9], PLID[3], CDS-CZSL[14], and Trokia[7]. The results are presented in Tab. 2. In the closed-world setting, our VP-CMJL achieves SOTA performance across all three datasets on nearly all metrics. Specifically, VP-CMJL improves the HM by $+5.5\%$, $+3.9\%$, and $+1.1\%$, and the AUC by $+3.9\%$, $+6.2\%$, and $+1.2\%$, respectively, across the three datasets, demonstrating the effectiveness of the proposed method. In the open-world setting, VP-CMJL significantly outperforms competing methods on the UT-Zappos and C-GQA datasets, with HM improvements of $+6.7\%$ and $+4.6\%$, and AUC improvements of $+8.4\%$ and $+1.9\%$, respectively. For the MIT-States dataset containing substantial label noise [2], the baseline method CDS-CZSL adopts dedicated pruning techniques specifically designed for open-world settings. Our method achieves competitive performance without relying on these task-specific optimizations.
4.3 Ablation Study
@@ 5 Conclusion
This work introduces the Visual Proxy and Cross-Modal Joint Learning strategy to address the challenges of Compositional Zero-Shot Learning (CZSL). By tackling the modality gap and the lack of fine-grained features, our approach introduces visual proxies for the first time and proposes a cross-modal joint learning strategy that enhances modality alignment and improves the discrimination of similar compositions, boosting generalization in CZSL tasks. Extensive experiments demonstrate that VP-CMJL outperforms existing methods in both closed-world and open-world settings, achieving state-of-the-art results. This work presents a promising direction for enhancing the generalization and discriminative power of compositional zero-shot learning by combining the strengths of textual and visual modalities.

### v4

@@ Abstract
Compositional Zero-Shot Learning (CZSL) aims to recognize novel attribute-object compositions by leveraging knowledge from seen compositions. Current methods align textual prototypes with visual features via Vision-Language Models (VLMs), but suffer from two limitations: (1) modality gaps hinder the discrimination of semantically similar pairs, and (2) single-modal textual prototypes lack fine-grained visual cues. In this paper, we introduce Visual Proxy Learning, a method that reduces modality gaps and enhances compositional generalization. We initialize visual proxies for attributes, objects, and their compositions using text representations and optimize the visual space to capture fine-grained cues, improving visual representations. Additionally, we propose Cross-Modal Joint Learning (CMJL), which imposes cross-modal constraints between the text-image and fine-grained visual spaces, improving generalization for unseen compositions and discriminating similar pairs. Experiments show state-of-the-art performance in closed-world scenarios and competitive results in open-world settings across four CZSL benchmarks, demonstrating the effectiveness of our approach in compositional generalization. The code will be available at https://github.com/codefish12-09/VP_CMJL.
@@ 1 Introduction
Figure 1: (a) The modality gap and coarse granularity of textual prototypes blur inter-class boundaries, leading to misclassification of similar samples (e.g., slightly red unripe apples). (b) The visual proxy improves modality alignment, integrates fine-grained image features, and makes the visual space more compact and separable, ensuring correct classification of similar samples.
Thanks to the advancements in large pre-trained vision-language models like CLIP [29], recent research has made significant progress by leveraging CLIP and exploring various approaches to enhance its capabilities. These approaches include designing task-specific prompts [24, 34, 35, 19, 3], decomposing compositional text features [18], and developing modal fusion methods [7] to narrow the modality gap. Such efforts have improved the alignment between compositional labels and image features, enabling CLIP to perform effectively in CZSL tasks. However, several challenges and inherent problems in the CZSL task still deserve attention.
Persistent Modality Gap. Despite efforts to narrow the modality gap between vision and language in existing methods, it persists as a fundamental challenge in CZSL. While prior work [17] demonstrates that the gap can be partially reduced, complete elimination remains unattainable, particularly under the current task setup where classification depends critically on top-1 cross-modal retrieval. In such scenarios, the distance between true cross-modal pairs may exceed that of false pairs, leading to misclassifications directly attributable to the modality gap (see Fig. 1(a)). This problem is further amplified by the fine-grained nature of compositional semantics, where subtle visual distinctions—such as texture variations between ’ripe’ and ’unripe’ apples—are semantically critical yet notoriously difficult to align with their textual counterparts.
Limited Discriminability of Textual Prototypes. Existing VLM-based CZSL methods learn textual prototypes to address label scarcity, yet these prototypes inherently lack the fine-grained visual details crucial for distinguishing similar compositions. While each visual category contains diverse image instances (e.g., apples in varying shapes or lighting conditions), textual prototypes are derived from a single compositional label, resulting in a semantic-visual asymmetry. This mismatch prevents prototypes from capturing subtle distinctions (e.g., ’ripe’ vs. ’unripe’ apples), leading to ambiguous decision boundaries in the visual space (Fig. 1(a)). As shown in prior studies [27], relying solely on textual prototypes fails to approach the optimal performance for visual tasks.
@@ 2 Related Work
Compositional Zero-Shot Learning. Previous methods for CZSL can be categorized into two strategies: (1) Uniform feature representation learning, which establishes relationships between compositions and primitives (attributes and objects), embedding both images and compositions into a shared space for unseen composition prediction using a single classifier [26, 22, 21, 1, 14]. (2) Multi-branch feature learning, which employs parallel discriminative modules for attributes, objects, and compositions, emphasizing visual feature discriminability [23, 11, 13, 32, 31, 16]. Recently, with the advancement of pre-trained Vision-Language Models (VLMs) and their strong generalization abilities, VLMs have been increasingly applied to CZSL [24, 34, 19, 18]. For instance, Nayak et al. [24] used soft prompts to treat attributes and objects as learnable tokens, while Xu et al. [35] constructed a graph [33] of object-attribute compositions. Lu et al. [18] decomposed states and objects in language features and integrated them with image features. Huang et al. [7] created three recognition branches to model attributes, objects, and compositions, aligning branch-specific prompt representations with visual features. Li et al. [15] proposed a context-based, diversity-driven specificity learning framework, considering attribute specificity levels.
@@ 3.1 Problem Formulation
In CZSL, compositional labels consist of attributes and objects. Attributes are denoted as $a\in\mathcal{A}$, objects as $o\in\mathcal{O}$, and compositions as $c\in\mathcal{C}$, where $\mathcal{C}=\mathcal{A}\times\mathcal{O}$. The compositional label space is divided into seen compositions $\mathcal{C}_{s}$ and unseen compositions $\mathcal{C}_{u}$, with $\mathcal{C}_{s}\cap\mathcal{C}_{u}=\varnothing$. The training set is defined as $\mathcal{T}=\{{(x,c)|x\in\mathcal{X},c\in\mathcal{C}_{s}\}}$, where $\mathcal{X}$ represents the image space. During testing, the model must classify images from both seen and unseen compositions. CZSL can be categorized into closed-world and open-world settings based on the treatment of unseen compositions [20]. In the closed-world setting, unseen compositions $\mathcal{C}_{u}$ are provided as prior knowledge, and the prediction space is $\mathcal{C}_{pred}=\mathcal{C}_{s}\cup\mathcal{C}_{u}$. In contrast, in the open-world setting, unseen compositions include all possible attribute-object compositions, and the prediction space is $\mathcal{C}_{pred}=\mathcal{A}\times\mathcal{O}$.
@@ 4.1 Experiments Setting
Table 2: Closed-world experimental results on the VAW-CZSL dataset, with the best performance in bold.
@@
4.2 Main Results
We comprehensively evaluate our model under both closed-world and open-world settings, comparing it with recent CLIP-based CZSL methods from the past two years, including: CLIP[29], CoOp[37], CSP[24], DFSP[18], DLM[6], ProLT[9], PLID[3], CDS-CZSL[15], Trokia[7] and IMAX[10]. The results are presented in Tab. 1. In the closed-world setting, our VP-CMJL achieves SOTA performance across all three datasets on nearly all metrics. Specifically, VP-CMJL improves the HM by $5.5\%$, $3.9\%$, and $1.1\%$, and the AUC by $3.9\%$, $6.2\%$, and $1.2\%$, respectively, across the three datasets, demonstrating the effectiveness of the proposed method. In the open-world setting, VP-CMJL significantly outperforms competing methods on the UT-Zappos and C-GQA datasets, with HM improvements of $6.7\%$ and $4.6\%$, and AUC improvements of $8.4\%$ and $1.9\%$, respectively. For the MIT-States dataset containing substantial label noise [2], the baseline method CDS-CZSL adopts dedicated pruning techniques specifically designed for open-world settings. Our method achieves competitive performance without relying on these task-specific optimizations.
4.3 Ablation Study
Table 3: Ablation experiment results on dual-modal center learning. ’TP’/’VP’ refers to text prototypes and visual proxies.
@@ 5 Conclusion
This work introduces the Visual Proxy and Cross-Modal Joint Learning strategy to address the challenges of Compositional Zero-Shot Learning (CZSL). By tackling the modality gap and the lack of fine-grained cues, our method introduces visual proxies for the first time and proposes a cross-modal joint learning strategy that enhances modality alignment and improves the discrimination of similar compositions, boosting generalization in CZSL tasks. Extensive experiments demonstrate that VP-CMJL outperforms existing methods in both closed-world and open-world settings, achieving state-of-the-art results. This work presents a promising direction for enhancing the generalization and discriminative power of compositional zero-shot learning by combining the strengths of textual and visual modalities.
Acknowledgements
This work is supported by the National Key R&D Program of China (No. 2023YFC3805203), National Natural Science Foundation of China (No. 62402354), and the Haikou Science and Technology Plan Project (2023-053).